Solve |x − 2| ≤ 5.
Answer by rothauserc(4718)

there are two cases to consider
:
1) x - 2 < or = 5
:
x < or = 7
:
2) -(x - 2) < or = 5
:
-x + 2 < or = 5
:
-x < or = 3
:
x > or = -3
:
**************************************
x < or = 7, x > or = -3
:
interval notation is [-3, 7]
**************************************
: